Castle Keep - Newcastle / Locations / News & Updates / United Kingdom

Castle Keep – Investigation Report


9 Sept. 2005
Conducted for Private Group

Chris Carnegie, Cindy Nunn, Colin Nunn


E.M.F. 0 throughout castle
Temperature: 15c throughout castle interior

Weather Conditions: dark and raining heavily

21:15 Compact disc player set up in the main entrance in the Great Hall. Scottish bagpipe music was played. This music was chosen as King Edward Longshanks is reported to haunt the keep and this would make him ‘appear’.

21:25 Motion detector placed in King’s Chamber.

21:30 Beam barrier placed in Gallery.

21:32 Beam barrier placed in Garrison Room.

21:37 Motion Detector placed in Chapel. £1 coin placed in chapel on piece of paper on top of the font.

21:42 Chess set placed in Great Hall. Both ‘Queens’ are placed on the wrong colour.

Whilst conducting the base readings the following observations occurred;

21:09 Chris and Colin both saw orbs in the Garrison Room with the naked eye.

21:20 Colin picks up on ‘man, long straight nose, white beard, long flowing blue gown’ in the passageway between the museum and the stairs.

21:25 Two dogs seen by Colin in the King’s Chamber, also ‘woman looking like ‘nursey’ from Blackadder, bubbly, giggly personality, not well-to-do but not a pauper’

21:26 Motion detector placed in the King’s Chamber was activated. Nobody was in the room at the time.

21:30 Colin sensed three males in the Gallery area ‘two hang around together, third is a loner, they aren’t amused’ Colin was disorientated.

The areas of the Keep investigated were Gallery, Garrison Room and Chapel, Kings Chamber and Main Hall and the Queen’s Chamber and anteroom.

Due to the employment of the rest of the investigators involved, API has decided it would be unwise to use the names of participants without their prior consent. In cases where consent has been given, only first names will be used. This is for the personal safety of the Police Officers involved and their families.


TEAM 1: Chris Carnegie, Cindy Nunn, Sharon and six others

VIGIL 1: Queen’s Chamber and anteroom

EQUIPMENT USED: digital cameras, 0lux video cameras, E.M.F. meters, temperature gun, thermometers, humidity meters, dowsing rods, pendulum


Time Witness Occurrence _

00:01 All A ‘tapping’ noise was heard coming from the wall.

00:03 All The same ‘tapping’ noise was heard.

00:06 All One of the E.M.F. meters starting ‘going crazy’ around one of
the participants.

00:10 All Both E.M.F. meters started playing up, one was showing a
constant high reading, the other was erratic. No wiring in the


00:19 ‘guest’ Saw a figure ‘small in height’ on the other side of the barriers
E.M.F. meter went off the scale at this point, room temp 27c

00:40 All A tapping noise was heard at the window on the landing, when
asked to repeat, it was heard again. Temp 30c on landing, 26c
in Queen’s Chamber.

00:41 All Both E.M.F. meters ‘off scale’ around Sharon, no results
around anyone else.

00:43 ‘guest’ ‘Feels’ a little girl next to her.

At this point the guests split into two groups, one in the Queen’s Chamber landing and one on the stairs. I stood at the top of the stairs to keep a record of what was happening. The rest of the notes for this vigil will be in two parts. The first from the Queen’s Chamber landing and the second from the stairs.
Part 1:
00:44 ‘guest’ Felt ‘something’ touch her foot

00:45 ‘guest’ Heard a small child laughing.

Part 2:
00:45 Sharon, 4 ‘guests’ A tapping noise was once again heard on the landing window.
00:47 ‘guest’ Felt something ‘tickle’ his ear.

END TIME: 00:51

TEAM 1: Chris Carnegie, Cindy Nunn, Sharon, Paul McDonald (Keep staff) and six others

VIGIL 2: Garrison Room and Chapel

EQUIPMENT USED: digital cameras, 0lux video cameras, E.M.F. meters, temperature gun, thermometers, humidity meters, dowsing rods, pendulum


Time Witness Occurrence _

01:40 All A motion detector was triggered and sounded. Upon closer
inspection it was found to be turned off. At this point Chris
switched it on.

01:41 All The motion detector was triggered again. No investigators
were in the vicinity. This was accompanied by strong E.M.F.
readings and a fall in temp from 22c to 15c in that area.

01:42 All Same as above but temp remained at 16.5c.

01:43- The scene described above kept happening, but not when any
01:46 All investigators were watching. It only seemed to happen when
we either turned away or ignored it. Nobody stood in the area
covered by the detector. Strangely it wasn’t activated after
had left the Chapel.


01:55 All Smell of damp in the Condemned Cell. E.M.F. meters both
erratic in this area.

02:08 All Paul seemed to be ‘on the verge of possession’ and sat
laughing uncontrollably.

02:10 All The team formed a circle in the middle of the floor, hoping
this would generate ‘paranormal’ activity.

02:15 Several ‘guests’ A cold draught was felt by several people in the circle, the
temperature was no different to anywhere else in the room,

02:25 Cindy Cindy felt as if something tapped her on her right shoulder.

02:25 All Scraping noise heard and the temperature suddenly dropped to

02:40 All A smell of stale tobacco was smelled by all team members. It
seemed to swirl around the room before disappearing.

02:46 All All team members started to complain about the cold, it had
fell to 14c in the room. The temp rose to 19c as the team left.

END TIME: 02:47

As several of the guests had left during this break and several factors outside of API’s control, it was decided to join the two teams and split the vigil between the Gallery and the Great Hall.

TEAM 1: Chris Carnegie, Cindy Nunn, Colin Nunn, Ivan, Sharon, Paul McDonald (Keep staff) and seven others

VIGIL 3: Gallery and Great Hall

EQUIPMENT USED: digital cameras, 0lux video cameras, E.M.F. meters, temperature guns, thermometers, humidity meters, dowsing rods, pendulum


Time Witness Occurrence _

03:30 All The team made a circle in the gallery and started breathing
exercises prior to having a séance.

03:34- All Colin and Cindy calling for spirits to come forward. One
03:50 camera turned off by itself, no other results.


03:55 Colin took ill at this point and had to be helped back down to the base room by Cindy. Chris stayed with the team.

04:20 Due to certain considerations, Chris conversed with Paul and it was decided to bring the invo to an end.

There are no Vigil reports for Team 2. The notebook containing these was not handed in.

This slideshow requires JavaScript.

Leave a Reply

Fill in your details below or click an icon to log in: Logo

You are commenting using your account. Log Out /  Change )

Google photo

You are commenting using your Google account. Log Out /  Change )

Twitter picture

You are commenting using your Twitter account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s